1) find all value of i^i, and show that they are all real

2) Find all values of log(-1-i)

3) find a) the cube roots of -1

b) the sixth root of i

c) the cube roots of 1-i

4) Find (d/dz) i^z
Answer #1

1.

а».fuf.»), nez i = elogr = eilogi-ei(logr+i@+2»)-JUgl+ǐは2m -μ+log r 0, (π<es..x=z.cosEy=rane =e(Mg.a.)=@ta»).neZ i = e (2 )